OpenStudy (anonymous):

Is the relationship shown by the data linear? If so, model the data with an equation. x y -7 5 -5 9 -3 13 -1 17

OpenStudy (anonymous):

A. The relationship is linear; y – 5 = 2(x + 7). B. The relationship is not linear. C. The relationship is linear; y – 5 = -1/2(x + 7). D. The relationship is linear; y + 7 = 1/2(x – 5).

OpenStudy (anonymous):

@team_cuban please help i will give medal

OpenStudy (anonymous):

@Linchpin

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

@team_cuban i have no idea im completly comfused and i need it solved

OpenStudy (anonymous):

I do not hand out answers but I can help you.

OpenStudy (anonymous):

even if i fan and medal ive been doing math for 6 hours straight

OpenStudy (anonymous):

Even if you fan and medal I still do not just hand out free answers. That would be taking a bribe and breaking COC now I can give you an easy way. Now for A, C, and D... plug in -7 as the x and 5 as the y and tell me what you get. For each, also show your work.

OpenStudy (anonymous):

A. 0=7 B.0=10.5 c.12=8.5 i simply just pluged them in and did the simple addition and division and multiplication

OpenStudy (anonymous):

Okay a little mistake on A. What is 2(-7+7)

OpenStudy (anonymous):

o its 0

OpenStudy (anonymous):

Okay so if 0=0 what do you suppose your answer is?

OpenStudy (anonymous):

not linear?

OpenStudy (anonymous):

the relationship in not linear or is the answer A?

OpenStudy (anonymous):

@team_cuban

OpenStudy (anonymous):

Correct it is A